OpenStudy (anonymous):

Find the sum of the geometric series with the first and last terms as given: a= 4, t6 = 972, r = 3

OpenStudy (anonymous):

Also, if you can: what is the sum of the first four terms of a geomtric sequence whose second term is 10 and whose fifth term is -1.25

OpenStudy (lgbasallote):

hmmm S_n = a1(1-r^n)/1-r we have a1 which is 4...r which is 3...and n which is 6 (from t6) so the only thing left to do is substitute....

OpenStudy (lgbasallote):

as for your second question you're given a_2 = 10 and T_5 = -1.25 T_5 = (a_2)r)r)r) T_5 = A_2 r^3 -1.25 = 10r^3 divide both sides by 10... -0.125 = r^3 cube root both sides... r = -0.5 can you take it from here??

OpenStudy (anonymous):

Im a lil confused, how do i get the terms?